Please Solve this rational equation:
x^3/16=x^6/2

Answer by mangopeeler07(462)

x^3/16=x^6/2

Cross multiply
16x^6=2x^3

Divide both sides by 2
8x^6=x^3

Divide both sides by x^3
8x^3=1

Divide both sides by 8
x^3=1/8

Cube root of 1/8= 1 over cube root of 8.
x=1/2